OpenStudy (anonymous):

Please HELP!!!! Need to show ALL work. If x=-5 is a root x^3+8x^2+16x+5, find the remaining roots.

OpenStudy (anonymous):

Plug -5 into the problem where x is and solve: (-5)^3+8(-5)^2+16(-5)+5 =-125+200-80+5 =0

OpenStudy (mathmale):

Know synthetic division? If so, that's a great tool for tackling this problem. Your job is to determine the remaining roots of the given polynomial.